# MAnycast Reloaded: a Tool for an Open, Fast, Responsible and Efficient Daily Anycast Census

### v1

@@ 4.1.1. Components
The tool consists of three components:
CLI – Takes a measurement as command-line input, and instructs the Orchestrator to perform that measurement.
Orchestrator – Central controller of MAnycastR.
Worker – Deployed at the anycast sites and receives instructions from the Orchestrator to probe the Internet.
Our design minimizes the burden on Workers by offloading computation to the Orchestrator whenever possible, allowing us to operate the census at a low cost. The separate CLI component allows measurements to be started from multiple environments (e.g., locally).
4.1.2. Measurement process
When starting a measurement, the CLI takes a measurement definition as input and forwards this to the Orchestrator. The Orchestrator then instructs all Workers that a measurement is starting, including a definition of the measurement such that Workers know what probes to send and listen for. The Workers then start listening. Next, the Orchestrator streams a list of IPs – at the CLI-defined probing rate – to each Worker. Workers send out probes as they receive hitlist targets from the Orchestrator. Each hitlist target receives a single probe from each Worker, the Orchestrator orchestrates it such that these probes are scheduled within a configurable offset after each other. When performing an ICMP-based measurement at an offset of 1 second, this means that from the target’s perspective they receive a set of probes from the same source address one second after each other, mimicking a regular ping sequence. We encode information regarding the sending Worker ID and the transmit time in fields that are echoed in responses from targets. For ICMP this is achieved using the ICMP payload, for DNS we encode information in the domain name of the request, and for TCP we use the acknowledgement number.
If a probed target is responsive, it sends back a response for each received probe, all these responses are routed to the nearest Worker, except for the case of the probed target being anycast as explained in §2. The Workers capture responses to probes and ensure they belong to the ongoing measurement by checking the information encoded in the probe that is echoed by the target. It then creates a result consisting of information encoded in the original probe and information such as the receive time and the receiving Worker ID. Results are streamed back to the Orchestrator, which receives streams from all Workers and in turn streams it to the CLI. At the CLI, results are stored as a single file aggregating received responses from all Workers.
@@ 4.1.3. Features
Unicast probingWorkers can perform a latency measurement using their unicast addresses (IPv4 and IPv6).
  • Latency-based measurements allow for accurate GCD-based anycast detection (R1).
Synchronized probing – A key limitation of MAnycast2 is that it performs measurements sequentially from individual VPs, leading to a significant time interval between outgoing probes to the same target. To address this limitation we use synchronized probing, where the Orchestrator instructs Workers to probe in parallel, with configurable offsets.
  • Route flips in-between probes are less likely to occur when probing in parallel, since each target receives probes in a short timeframe, reducing the number of FPs (R1, R2).
@@
  • The measurement is performed at all Workers in parallel, rather than for each Worker individually (R5, R11).
@@
Additional protocols – MAnycastR supports ICMP, TCP, and UDP/DNS probing. TCP probing uses SYN/ACK packets to high port numbers, for which we receive RST packets. Since anycast is widely used for DNS infrastructures, we implemented DNS probing by sending out A record queries. We also support TXT record queries for the CHAOS class.
  • Coverage is increased by supporting service-aware UDP/ DNS and TCP probing (R4).
@@
Failure awareness – When a Worker disconnects during a measurement, the Orchestrator completes the measurement with the remaining Workers. Disconnecting the CLI can be used to cancel incorrect measurements.
@@
  • Misconfigured measurements can be aborted, such that no unnecessary probes are sent out (R3).
Buffering and aggregation – At the start of a measurement, the CLI sends the hitlist to the Orchestrator, which buffers it and streams it to the Workers at the CLI-defined probing rate. Furthermore, results captured at Workers are forwarded to the Orchestrator for aggregation.
@@
Containerization – MAnycastR supports Docker containerization, with an image size of $\sim$130 MB.
@@
  • Docker allows for easy, OS-independent deployment. Furthermore, a small image size results in easy distribution and low storage requirements (R7, R9, R10).
Secure inter-component communication – MAnycastR can be deployed using an Orchestrator certificate and private key, requiring public keys at the Workers.
  • Components are authenticated and communication between components is encrypted (R8).
All source code is made available under a permissive license (Mozilla, 2025) with documentation to satisfy R9. The final measurement system allows for a large range of measurements, e.g, anycast-based, GCD, and, e.g., anycast catchment measurements (de Vries et al., 2017). Furthermore, it makes a daily anycast census feasible both in terms of probing time and probing cost.
Platform Anycast/unicast # of VPs Limitations
MAnycastR production Both 32 Few VPs for GCD
Archipelago (Claffy et al., 2009)
Unicast only
Up to 180 (IPv4)
Up to 100 (IPv6)
Suitable for GCD only
Fewer nodes IPv6/TCP compatible
Table 1. Measurement platforms used in this work.
4.2. Measurement pipeline
In this section we describe the measurement platforms we use and our pipeline as implemented on these platforms.
4.2.1. Measurement platforms
Our pipeline makes use of two measurement platforms: MAnycastR production and CAIDA’s Archipelago (Ark) measurement platform (Claffy et al., 2009).
MAnycastR production is our anycast deployment that is deployed on Vultr’s cloud and makes use of all its 32 sites, located in 19 countries on 6 continents (Vultr, 2024). We announce a /24 IPv4 prefix and a /48 IPv6 prefix from all sites. For latency-based GCD measurements we need a large number of geographically distributed VPs, for which we use the Ark platform. Ark has 180 VPs that allow for measurements using the scamper tool (Luckie, 2010). We chose Ark due to its reliability, well-documented VP locations, and good geographical coverage to allow for accurate longitudinal GCD measurements. Table 1 specifies the properties of both platforms.
Figure 3. Overview of the measurement pipeline.
4.2.2. Pipeline
Figure 3 shows the measurement pipeline for our daily anycast census. The MAnycastR system is deployed on the anycast deployment. Each anycast site runs a Worker instance, the Orchestrator instance is deployed on-premise at our institution. The CLI component used to trigger the daily census is also deployed on-premise. This setup performs anycast-based measurements for each supported protocol (ICMP, TCP, DNS) for both IPv4 and IPv6. We perform GCD measurements using ICMP and TCP, we do not use DNS due to the possible jitter introduced by DNS request processing by the target.
@@ 4.2.2. Pipeline
Given the low probing cost and feasibility, we perform the anycast-based measurements using the anycast deployment toward full hitlists (IPv4 and IPv6). This measurement leads to a set of candidate anycast prefixes, which we call anycast targets (AT). We extend these ATs to include anycast prefixes found with GCD in prior measurements (purple arrow) allowing our census to also cover rare cases where the anycast-based methodology fails to detect anycast (i.e., FNs of MAnycast2). The initial set of prefixes, fed back into the measurement, is found using large-scale GCD measurements and operator ground truth, which we discuss later.
@@
Next, using the Ark platform we perform GCD measurements towards the aforementioned candidate ‘anycast targets’. This step allows us to enumerate and geolocate anycast sites using iGreedy’s analysis approach. By performing the GCD measurement only on ‘anycast targets’ we significantly reduce the probing cost by two orders of magnitude, from millions to tens-of-thousands. This reduction is necessary as measuring the entire hitlist with GCD is impossible for a daily census at responsible probing rates that do not overburden the Ark platform nor the Internet.
We find that Ark provides sufficient geographical coverage for GCD. Whilst we acknowledge that more VPs lead to higher enumeration of anycast sites, we believe the burden and probing cost of performing GCD using even larger scale measurement platforms outweighs the value. Furthermore, fully enumerating large anycast deployments is infeasible as backed by previous work (Fan et al., 2011). We later confirm this by sharing results of GCD measurements using 481 geographically distributed RIPE Atlas probes (§5.2).
4.2.3. Probing strategy
Our ping targets come from ISI’s IPv4 hitlist (Fan and Heidemann, 2010), which ranks ping-responsive addresses per /24 prefix. For DNSv4 we merge the ISI hitlist with a set of authoritative name server IPs obtained from the OpenINTEL project (van Rijswijk-Deij et al., 2016), and we prefer name server IPs as representative addresses for a specific /24 over the ISI hitlist. We adopt this approach to maximize chances of probing an active DNS server. For IPv6 we use the IPv6 hitlist from TU Munich (TUM) (Gasser et al., 2018) combined with AAAA DNS records obtained from the OpenINTEL project. Our IPv4 hitlist totals 5.9 M targets of which $\sim$4.0 M responsive (depending on churn), and for IPv6 we have 2.5 M targets of which $\sim$1.7 M responsive. We regularly update the hitlists to improve coverage of the census. In particular, we observe the IPv6 hitlist sources to show large increases in coverage of IPv6 address space over time.
We probe at a /24-IPv4-prefix and /48-IPv6-prefix granularity as these are likely to be the smallest prefix size to be propagated by BGP. Whilst this assumption works well in most of the cases, Schomp et al. showed that there may be rare cases of more specific partitioning (Schomp and Al-Dalky, 2020). For example, hypergiants announcing a /24 prefix at their PoPs may assign part of a prefix to a replicated service, whilst utilizing all other addresses in that prefix as unicast for individual servers within their private backbone. While we acknowledge this limitation, we continue to probe at a /24 prefix granularity to keep our scanning rate low for responsible scanning. We address and quantify this limitation in §5.6.
4.2.4. Daily output
@@
Our daily census is finally constructed as follows. For each prefix where either the anycast-based or GCD methodology detects anycast, we publish the results to our public Git repository1. The census includes information found using both methodologies for all protocols. Furthermore, we add the estimated number of anycast sites found by each measurement, and for the latency-based measurements we include expected geolocations generated using iGreedy’s city population-based geo-detection algorithm. This data can be used by the community to evaluate anycast deployments with different levels of confidence.
4.3. Ethics
Our ethics review board does not require a specific ethics approval for measurements conducted according to community best practices. In this work, we follow these practices by ensuring that the address space of our measurement infrastructure has correct pointers to abuse contacts and references to a page explaining the goal of the measurements, with clear instruction for opting out. The analysis conducted in this paper was also helpful in establishing a significantly lower and manageable probing rate. This allows us to perform a responsible daily census that minimises performance impact on the targeted infrastructure.
5. Validation
In this section we validate the performance of MAnycastR.
@@ 5. Validation
We focus on: (§5.1) accuracy and precision, (§5.2) accuracy of site enumeration, (§5.3) the benefits of increased protocol coverage, (§5.4) replicability in other deployments, (§5.5) accuracy at low probing rates, (§5.6) prefix partitioning, (§5.7) comparison against other censuses and (§5.8) ground truth.
@@ 5.1.1. Large-scale GCD measurement
As discussed in §2, GCD is highly accurate in detecting anycast. However, the probing cost for GCD measurements on the entire hitlists is too high for frequent measurements. Because of the accuracy benefits of such a measurement, we conducted GCD measurements on the entirety of both hitlists in February and December of 2024. We analyze the latter measurement run in this section. In this case we used 227 Ark nodes for ICMPv4 and 118 for ICMPv6. The number of Ark nodes is higher than our daily census, as it made use of the Ark development environment, which has more VPs. To limit impact, we performed the run at a low probing rate over a period of 22 days for IPv4 and 9 days for IPv6. For clarity, we refer to this validation measurement as GCDArk  [^3] Note: the MAnycast2 paper also discusses a GCD measurement, limited to a 2% sample of prefixes to save on probing cost (§4.4 of (Sommese et al., 2020)).. This measurement provides a lower bound of the anycast deployments on the Internet, since using the GCD approach tends to underestimate the number of anycast deployments (§2.1). The results of this validation run provide insight into the number of FNs of the anycast-based measurement and will be used as a baseline for further validation. Table 2 shows the results of this measurement, compared to the candidate anycast targets found using the anycast-based approach of MAnycastR (red list marked “AT” in Figure 3).
[…]
ICMPv4 25,396 13,692 13,168 524 (3.8%) 12,228
ICMPv6 6,315 6,221 6,006 215 (3.5%) 94
Table 2. Comparing anycast prefixes found using the anycast-based measurement to GCDArk.
The GCDArk ICMPv4 measurement finds 13,692 anycast prefixes, of which 13,168 are also found by the anycast-based measurement in the MAnycastR pipeline. Ergo, the pipeline misses 524 prefixes. Of these 524 prefixes, 365 are covered by our public census as they were added to the AT list after the GCDArk scan in February. As for accuracy, the anycast-based approach finds 25,396 anycast targets, of which 12,228 were classified as unicast by GCDArk. While these are mostly FPs of the anycast-based approach, ground truth shows these also contain TPs that we discuss later. For IPv6 we find 6,221 anycasted /48s using GCDArk and 6,315 using MAnycastR and an intersection of 6,006 /48s.
This analysis shows there is value in doing large-scale latency measurements, as this covers prefixes missed by the anycast-based approach. However, the drawbacks in terms of probing cost, timing, and burdening of infrastructure outweighs the gain in visibility for daily measurements. To balance the benefits, we repeat large-scale GCD measurements bi-annually and feed the additional prefixes found into our AT list so they are covered in subsequent daily census results.
@@ 5.1.2. RIPE Atlas GCD measurement campaign
To assess whether the daily pipeline provides sufficient geographical coverage for GCD measurements, we ran an extensive measurement campaign using 481 RIPE Atlas nodes. We selected nodes with stable uptime, filtered out nodes with false user-reported locations (by verifying with both MaxMind and IP2location databases (IP2Location, 2025; MaxMind, 2025)), and ensured that no two nodes are within 100 km of each other. We conducted the measurement towards the ICMPv4 ATs of 16 September, 2024, which consisted of 23,821 prefixes.
Comparing the results, we detect 12,186 prefixes as anycast using Ark in the MAnycastR GCD measurement and 12,202 prefixes using RIPE Atlas, with an intersection of 11,953 prefixes. RIPE Atlas missed 233 prefixes found only in the MAnycastR census, mostly due to probe measurement failures. With RIPE Atlas we find 248 prefixes not found using GCD in the census. Of these, 189 belong to Imperva, a CDN that offers DDoS protection. We suspect these to be cases of temporary anycast, as we discuss later. For the remaining 59 prefixes we see cases of a single RIPE Atlas node reporting low RTT, which we suspect to be ASes blocking content and regional anycast deployments where MAnycastR lacks VPs.
Overall these results show that RIPE Atlas provides little additional coverage. Whilst we find the higher geographical coverage to increase enumeration capabilities, as we will discuss later, it incurs a significant increase in probing cost. Furthermore, even with increased spending quotas the measurement lasted three days. For these reasons we find RIPE Atlas to be unsuitable for the daily MAnycastR pipeline. We provide more detail in Appendix A

### v3

@@ 4.2.1. Components
Our anycast-based probing system consists of three components:
Orchestrator – Central controller for anycast probing, deployed on-premise at our institution.
CLI – Interface to create a measurement definition and instruct the Orchestrator to perform a specific measurement.
Worker – Deployed at the anycast sites to receive instructions from the Orchestrator to probe the Internet.
We minimize the burden on Workers by offloading all computation to the Orchestrator except for sending, receiving, and matching of packets. The separate CLI component allows measurements to be started from multiple environments. We deploy Workers on TANGLED (Bertholdo et al., 2021), our anycast testbed that is deployed on Vultr’s cloud and uses of all its 32 sites, located in 19 countries on 6 continents (Vultr, 2025b). We announce a /24 IPv4 prefix and a /48 IPv6 prefix from all sites. In §5.5.1 we motivate our decision to use Vultr.
4.2.2. Daily measurement process
When starting a measurement, the CLI takes a measurement definition as input and forwards this to the Orchestrator. The Orchestrator then instructs all Workers that a measurement is starting, including a definition of the measurement such that Workers know what probes to send and listen for. The Workers then start listening. Next, the Orchestrator streams a list of IPs – at the CLI-defined probing rate – to each Worker. Workers send out probes as they receive hitlist targets from the Orchestrator. Each hitlist target receives a single probe from each Worker, the Orchestrator orchestrates it such that these probes are scheduled within a configurable offset after each other. When performing an ICMP-based measurement at an offset of 1 second, a target will receive a set of probes from the same source address one second after each other, mimicking a regular ping sequence. We encode the sending Worker ID and the transmission time in fields that are echoed in responses from targets. For ICMP we use the ICMP payload, for DNS we encode information in the domain name of the request, and for TCP we use the acknowledgement number.
If a probed target is responsive, it sends back a response for each received probe and these responses are routed to the nearest Worker, except for the case of the probed target being anycast as explained in §2. The Workers capture responses to probes and ensure they belong to the ongoing measurement by checking the information encoded in the probe that is echoed by the target. The Worker then creates a result consisting of information encoded in the original probe and information such as the receive time and the receiving Worker ID. Results are streamed back to the Orchestrator that receives streams from all Workers and aggregates results in a single stream to the CLI. At the CLI, results are stored as a single file.
@@ 4.2.3. Features
Synchronized probing – A key limitation of the approach in (Sommese et al., 2020) is that it performed measurements sequentially from individual VPs, leading to a significant time interval between outgoing probes to the same target. To address this limitation, we use synchronized probing, where the Orchestrator instructs Workers to probe in parallel.
  • Route changes in-between probes are less likely to occur when probing in parallel, since each target receives probes in a short timeframe, reducing the number of FPs (R1).
  • Reducing FPs from route changes (variable between subsequent censuses) improves longitudinal result stability (R2).
@@
  • Measurements are performed by all Workers in parallel, rather than sequentially, thus improving scalability (R5).
@@
Additional protocols – Our system supports ICMP, TCP, and DNS over UDP probing. TCP probing uses SYN/ACK packets to high port numbers, for which we receive RST packets. Since anycast is widely used for DNS infrastructures, we implemented DNS probing with A and CHAOS TXT record queries.
  • Coverage is increased by supporting DNS and TCP probing (R4).
@@
Failure awareness – When a Worker disconnects during a measurement (e.g., due to outside factors like link failure) the measurement is completed by the Orchestrator using the remaining Workers. Disconnected Workers automatically reconnect.
@@
Buffering and aggregation – At the start of a measurement, the CLI sends the hitlist to the Orchestrator, which streams it to the Workers at the CLI-defined probing rate. Results captured at Workers are forwarded to the Orchestrator for aggregation.
@@
Deployment – Our anycast-based probing system can be deployed using a statically-linked binary (8 MB) on both Linux and FreeBSD operating systems or Docker (image size of $\sim$8 MB).
  • Docker allows for easy, OS-independent deployment (R6).
@@
  • A small image and binary size result in easy distribution and low storage requirements (R6).
The source code is written in Rust and is publicly available with a permissive license (Mozilla, 2025) with extensive documentation (Hendriks, 2025a).
4.3. Enumeration and Geolocation Pipeline
@@
Figure 3 shows the measurement pipeline for our daily anycast census. Given the low probing cost and feasibility, we perform the anycast-based measurements using TANGLED (§4.2) toward full hitlists (§4.1). This measurement leads to a set of candidate anycast prefixes, which we call anycast candidates (AC). We extend these ACs to include anycast prefixes found with GCD in prior measurements (purple arrow) allowing our census to also cover cases where the anycast-based methodology fails to detect anycast (i.e., FNs of MAnycast2). The initial set of prefixes, fed back into the measurement, is found using large-scale GCD measurements and operator ground truth, which we discuss later (§5.1.1).
We perform GCD measurements towards the ACs. For latency-based GCD measurements we need a large number of geographically distributed VPs. The first two months of the census, GCD measurements were conducted using TANGLED (32 VPs). Since Jun. ’24, we perform these measurements using CAIDA’s Ark (Claffy et al., 2009), which initially increased the number of VPs to $\sim$160 for ICMPv4 and $\sim$90 for ICMPv6. During the census, Ark has grown substantially as we use $\sim$250 ICMPv4 and $\sim$150 ICMPv6 VPs in Sep. ’25.
@@
We chose Ark due to its reliability, well-documented VP locations, and geographical coverage. Furthermore, Ark provides a Python API allowing for efficient collection of results (Luckie et al., 2025). GCD measurements allow us to enumerate and geolocate anycast sites using iGreedy’s analysis approach, for which we make an improved version publicly available that severely reduces processing time (Hendriks, 2025b). By performing the GCD measurement only on ACs we significantly reduce the probing cost by two orders of magnitude, from millions to tens-of-thousands. This reduction is necessary as measuring the entire hitlist with GCD is impossible for a daily census at responsible probing rates that do not overburden the Ark platform.
The component that performs anycast-based measurements used ICMP, TCP and DNS for both IPv4 and IPv6. We performed GCD measurements using ICMP and TCP. We did not use DNS due to the possible jitter introduced by DNS request processing by the target that may inflate captured latency and affect the detection algorithm. We have a monitoring system in place that warns when an upload fails, when few VPs participate, or results deviate from the baseline.
4.4. Census Output
@@
After this step we have two groups of results that are published in the census. First, we have $\mathscr{G}$ which are ACs confirmed with the follow-up GCD measurement. Secondly, we use $\mathscr{M}$ to describe ACs that are not confirmed with GCD. Our daily census is finally constructed as follows. For each prefix where either the anycast-based ($\mathscr{M}$) or GCD methodology ($\mathscr{G}$) detects anycast, we publish the results to our public Git repository (git, 2025). The census includes information found using both methodologies for all protocols. This data can be used by the community to evaluate anycast deployments with different levels of confidence. Furthermore, we add the estimated number of anycast sites found by each measurement, and for the latency-based measurements we include expected geolocations generated using iGreedy’s city population-based geo-detection algorithm. We also include the number of VPs that participated, which is variable since Ark grows over time, as it impacts the accuracy of results (see §7 for more details).
5. Performance Assessment
Figure 4. Roadmap of analysis conducted, detailing the date at which it was performed, the section in which we discuss the results, and the associated figures & tables.
@@ 5. Performance Assessment
In this section we assess the performance of LACeS. We focus on: (§5.1) accuracy and precision, (§5.2) site enumeration, (§5.3) benefits of broader protocol coverage, (§5.4) replicability, (§5.5) accuracy at low probing rates, (§5.6) impact of routing topology, (§5.7) prefix partitioning and (§5.8) comparison against other censuses. As we assess a longitudinal dataset, we conduct analyses using data collected att different points of time, summarized in Figure 4.
@@ 5.1. Accuracy and precision
Protocol AC GCDLS $\textbf{AC}\cap\textbf{GCD\textsubscript{{LS}}}$ FNs (FNR%) $\neg\textbf{GCD\textsubscript{{LS}}}$
ICMPv4 30,609 13,514 12,698 (94.0%) 816 (6.0%) 17,911
ICMPv6 11,166 11,286 10,617 (94.1%) 669 (5.9%) 549
Table 1. Comparing anycast candidates (AC), excluding candidates from the feedback loop, to GCDLS.
@@ 5.1.1. Large-scale GCD measurement
As discussed in §2, GCD is highly accurate in detecting anycast. However, the probing cost for GCD measurements on entire hitlists is too high for frequent measurements. Because of the accuracy benefits of such a measurement, we perform periodic GCD measurements on the entirety of our IPv4 and IPv6 hitlists. To limit impact, we performed the run at a low probing rate of 100 packets per second over a period of several days. For clarity, we refer to this measurement as GCDLS  [^1] Note: the MAnycast2 paper also discusses a GCD measurement, limited to a 2% sample of prefixes to save on probing cost (§4.4 of (Sommese et al., 2020)).. This measurement provides a lower bound of the anycast deployments on the Internet, since using the GCD approach can underestimate the number of anycast deployments (§2.1). We conducted GCDLS measurements in Feb. ’24, Dec. ’24, and Aug. ’25 finding 13,684, 13,692, 13,514 /24s respectively. Table 1 shows the results for GCDLS from Aug. ’25, compared to the ACs found using the anycast-based approach of LACeS (“AC” in Figure 3).
The GCDLS ICMPv4 measurement inferred 13,514 anycast prefixes, of which 12,698 (94.0%) are also inferred by the anycast-based measurement in the LACeS pipeline. Ergo, the anycast-based approach missed 816 (6.0%) prefixes. These prefixes are included using the feedback loop, thus providing coverage of these anycast-based FNs. In §7 we discuss the longitudinal validity of the feedback loop.
As for accuracy, the anycast-based approach finds 30,609 ACs, of which 17,911 (58.5%) were classified as unicast by GCDLS. While these are mostly FPs of the anycast-based approach, ground truth shows these also contain TPs that we discuss in §6. For IPv6 we find 11,286 anycasted /48s using GCDLS and 11,166 using LACeS and an intersection of 10,617 /48s. This analysis shows there is value in doing large-scale latency measurements, as this covers prefixes missed by the anycast-based approach. However, there are drawbacks in terms of probing cost, timing, and infrastructure burden. To balance the benefits, we repeat large-scale GCD measurements periodically and feed the prefixes found into our AC list so they are covered in subsequent daily census results.
@@ 5.1.2. RIPE Atlas GCD measurement campaign
To assess whether the daily pipeline provides sufficient geographical coverage for GCD measurements, we ran an extensive measurement campaign using 481 RIPE Atlas nodes in Sep. ’24. We selected nodes with stable uptime, user-reported location consistent with both MaxMind (MaxMind, 2025) and IP2location (IP2Location, 2025) and ensured that no two nodes are within 100 km of each other. We conducted the measurement towards 23,821 ICMPv4 ACs found on the 16th of September, 2024.
Comparing the results, we detect 12,186 (51%) prefixes as anycast in the LACeS GCD measurement ($\mathscr{G}$), using 164 Ark VPs at the time, and 12,202 (51%) prefixes using RIPE Atlas, with an intersection of 11,953 prefixes. RIPE Atlas missed 233 prefixes confirmed using GCD in the LACeS census, mostly due to probe measurement failures. With RIPE Atlas we find 248 prefixes not found using GCD in the census. Of these, 189 belong to Imperva, a CDN that offers DDoS protection. We suspect these to be cases of temporary anycast which we discuss in §7. For the remaining 59 prefixes we see cases of regional anycast in areas where LACeS lacks VPs.
Overall these results show that RIPE Atlas provides relatively small gains in identifying anycast prefixes, and in enumeration of sites as we show in §5.2. Yet, even with increased spending quotas the measurement lasted three days which is why we find it challenging to use RIPE Atlas in a daily LACeS pipeline. We provide more detail in Appendix B.